# The Art of Imitation: Learning Long-Horizon Manipulation Tasks from Few Demonstrations

### v1

@@ IV-E Time-Based Initialization
EM is not guaranteed to converge to the global optimum, but only to a local optimum [10]. Consequently, its initialization is of critical importance. $K$-means clustering is popular but does not guarantee a global optimum either [10]. We efficiently segment the trajectories along the time dimension using $K$-bins. Fig. 4 illustrates that complex task demonstrations are usually not aligned temporally. However, our skill segmentation proposed in Sec. IV-C aligns them, yielding a good initialization for EM. The combination of skill segmentation and time-based initialization enables the EM procedure to arrive at good local optima, while also achieving significantly faster convergence. Next, we discuss how we leverage it to efficiently select task parameters, too.
@@ IV-F Task-Parameterization
Previous work has usually assumed groundtruth access to relevant task-parameters [9, 8] or used infrared motion tracking devices [10], both of which restrict the applicability of the method. We therefore propose an approach to automatically select the relevant task parameters from RGB-D observations. Moreover, different skills in the same task require different information, e.g., in pick and place, the picking skill is often independent of the placing site. Modelling the full task using both task parameters is not only more computationally expensive but also leads to worse task performance. Hence, our parameter selection method has a synergistic effect with the skill segmentation proposed in Sec. IV-C. We offer a modular two-staged approach to candidate generation based on visual features and task-parameter selection based on statistical considerations. Both operate independently of each other, rendering them straightforward to substitute.
@@ IV-F1 Candidate Generation
Various methods exist for extracting task parameters from visual observations [23]. Due to their generalization capabilities, we build on recent advances in 3D scene keypoints [4] and leverage a DINO vision transformer [12]. We sample a set of reference descriptors from the embedding of the first image observation of task demonstrations, either manually in a GUI or automatically using SAM-generated object masks [34]. The associated keypoints constitute the set of candidate task parameters. We then localize them in each demonstration by establishing correspondences via the embedding distances [4, 12].
@@ IV-F2 Candidate Selection
In contrast, Alizadeh et al. fit a joint model to all candidates and the full task trajectory and calculate the relative precision during trajectory reconstruction [11]. Our per-skill single-frame approach scales to long-horizon tasks and large sets of candidate parameters while the temporal alignment eliminates the need for a reconstruction phase.
TABLE III: Time-driven vs. state-driven TAPAS-GMM, with and without disturbance of the rollout. Among the time-driven policies we compare different post-processing strategies Among the state-driven policies, we compare naive and factorized velocities. Trajectory length is mean and standard deviation.
@@
TABLE IV: Real world policy success rates.
@@ V Experimental Results
We first evaluate the efficacy of TAPAS-GMM for policy learning on a set of challenging manipulation tasks from the well established RLBench benchmark [13] and ablate the major components of our method. Afterward, we compare time-driven and state-driven policies on custom tasks implemented in ManiSkill2 [35]. Finally, we verify our results on a real robot and demonstrate our method’s generalization capabilities. Fig. 6 illustrates the evaluated RLBench tasks and real world tasks. Additional details for all the experiments are provided in Sec. S.4 of the supplementary material.
@@ V-A Time-Driven Policy Learning in Simulation
We compare against LSTM [33] and Diffusion Policy [32] baselines trained with either 100 or 5 demonstrations. In contrast, TAPAS-GMM is only trained on five demonstrations. To experimentally disentangle policy learning and representation learning, we use groundtruth object poses as inputs for the baselines and as candidate task parameters for TAPAS-GMM. We evaluate our proposed visual features in Sec. V-C and Sec. V-D. We further compare against 3D Diffuser Actor [31], which at the time of writing is the incumbent on RLBench. Note that it leverages a 3D scene representation instead of object poses. We evaluate all policies for 200 episodes. Tab. I shows that TAPAS-GMM matches or outperforms the SOTA across all tasks while requiring significantly fewer demonstrations. Moreover, in contrast to all baselines, TAPAS-GMM scales to more complex long-horizon tasks such as SweepToDustpan and PlaceCups, without loss in performance. Moreover, unlike the Diffusion Policy and LSTM, it excels when high precision is required, e.g., in PushButton.
@@
We compare the effectiveness of time-driven and state-driven GMMs in Tab. III. To simulate disturbances during policy rollout, we reset the end-effector pose back to the starting pose, around the object contact point in the trajectory. While state-based policies are harder to learn from few demonstrations, they outperform time-based policies when confronted with disturbances due to their time-invariance. Furthermore, our action-factorized policies consistently outperform the naive implementation.
@@ V-C Real-World Policy Learning
The policy success rates in Tab. IV are consistent with the results in our simulated experiments and demonstrate the core advantages of our approach: 1. strong generalization from few demonstrations, 2. handling of articulated objects, and 3. scaling to complex tasks and long horizons. Consistent with our simulated ablation experiment, TP-GMMs struggle with pose accuracy and gripper timing due to locally irrelevant task parameters and poor temporal alignment. Moreover, the deep learning baselines fail to generalize from five demonstrations and do not learn to solve any task instance. In contrast, TAPAS-GMM effortlessly scales to long-horizon tasks, without any loss in accuracy or generalization performance.
@@ V-D Visual Generalization and Skill Reuse
Prior work has extensively demonstrated the generalization capabilities of keypoints as visual features [4], which our policies inherit through the parameter generation proposed in Sec. IV-F1. We study the resulting zero-shot generalization capabilities of our policies on PickAndPlace and PourDrink by evaluating the policy with unseen objects instances, clutter objects, and task environments. We observe strong generalization performance across the board. We further provide qualitative results and discussions in the supplementary video.
Another advantage of TAPAS-GMM is that it not only solves one task but instead learns skill models, which can be recombined to solve novel tasks. In the supplementary video, we demonstrate how the segmented skills can be sequenced and reused in novel ways. In particular, the BakeBread task consists of taking the tray out of the oven and placing the bread on it. As an example, we invert and reorder these skills to construct a policy that takes the bread from the tray and inserts the tray into the oven.
@@ VI Conclusion
We present TAPAS-GMM, a method for learning complex manipulation tasks from only five demonstrations. By segmenting skills and automatically selecting relevant task parameters per skill, our method scales both to tasks with long horizons and many parameters. Through leveraging visual semantic features, our method further generalizes across object instances, environments, and clutter. Finally, our proposed velocity factorization makes time-invariant task-parameterized policies more effective.
Limitations: TAPAS-GMM is vulnerable to some of the core problems of imitation learning. First, it does not reason about the robot’s kinematic limits or object collisions. Second, it requires a decent coverage of the task space to generalize, which can be difficult to achieve on multimodal trajectory distributions, when given only five demonstrations. This shows in the PlaceCups task, where workspace limits cause different task instances to require vastly different trajectory shapes. Adding a prior, encoding workspace limits, and object collision probabilities, might alleviate this problem.

### v2

@@ IV-D Skill Sequencing and Skill Reuse
TABLE III: Training and inference times across RLBench tasks. Inference times are normalized by the prediction horizon of the policy.
[…]
5 Demos 100 Demos
TAPAS-GMM 0:00 – 0:01 00:02 – 00:11 000100
TP-GMM 0:00 – 0:03 00:05 – 02:00 000100
LSTM 0:40 – 1:00 02:00 – 04:30 007.000
Diffusion Policy 1:30 – 5:00 30:00 – 99:00 100.000
@@ IV-E Time-Based Initialization
EM is not guaranteed to converge to the global optimum, but only to a local optimum [10]. Consequently, its initialization is of critical importance. $K$-means clustering is popular but does not guarantee a global optimum either [10]. In contrast, we efficiently divide the trajectories along the time dimension into $K$ bins. Fig. 3 illustrates that complex task demonstrations are usually not aligned temporally. However, our skill segmentation proposed in Sec. IV-C aligns them and samples the segments to a common length so that $K$-bins yields a good initialization for EM. The combination of skill segmentation and time-based initialization enables the EM procedure to arrive at good local optima while also achieving significantly faster convergence. Next, we discuss how we leverage it to efficiently select task parameters, too.
@@ IV-F Task-Parameterization
Previous work has usually assumed groundtruth access to relevant task-parameters [9, 8] or used infrared motion tracking devices [10], both of which restrict the applicability of the method. Therefore, we propose an approach to automatically select the relevant task parameters from RGB-D observations. Moreover, different skills in the same task require different information, e.g., in pick and place, the picking skill is often independent of the placing site. Modeling the full task using both task parameters is not only more computationally expensive but also leads to worse task performance. Hence, our parameter selection method has a synergistic effect with the skill segmentation proposed in Sec. IV-C. We offer a modular two-staged approach to candidate generation based on visual features and task-parameter selection based on statistical considerations. Both operate independently of each other, rendering them straightforward to substitute.
@@ IV-F1 Candidate Generation
Various methods exist for extracting task parameters from visual observations [23]. Due to their generalization capabilities, we build on recent advances in 3D scene keypoints [4] and leverage a DINO vision transformer [12]. We sample a set of reference descriptors from the embedding of the first image observation of task demonstrations, either manually in a GUI or automatically using SAM-generated object masks [34]. The associated keypoints constitute the set of candidate task parameters. We then localize them in each demonstration by establishing correspondences via the embedding distances [4, 12]. In our experiments, we then set the orientation of all keypoint frames to the identity quaternion, which we found sufficient for solving all evaluated tasks. If the precise orientation of the keypoint frames should be required, it can be estimated using off-the-shelf methods [23].
@@ IV-F2 Candidate Selection
In contrast, Alizadeh et al. fit a joint model to all candidates and the full task trajectory and calculate the relative precision during trajectory reconstruction [11]. Our per-skill single-frame approach scales to long-horizon tasks and large sets of candidate parameters, while the temporal alignment eliminates the need for a reconstruction phase.
TABLE IV: Time-driven vs. state-driven TAPAS-GMM, with and without disturbance of the rollout. Among the time-driven policies we compare different post-processing strategies Among the state-driven policies, we compare naive and factorized velocities. Trajectory length is mean and standard deviation.
@@
TABLE V: Real world policy success rates.
@@ V Experimental Results
We first evaluate the efficacy of TAPAS-GMM for policy learning on a set of challenging manipulation tasks from the well established RLBench benchmark [13] and ablate the major components of our method. Afterward, we compare time-driven and state-driven policies on custom tasks implemented in ManiSkill2 [35]. Finally, we verify our results on a real robot and demonstrate our method’s generalization capabilities. Fig. 6 illustrates the evaluated RLBench tasks and real-world tasks. Additional details for all the experiments are provided in Sec. S.4 of the supplementary material.
@@ V-A Time-Driven Policy Learning in Simulation
We compare against LSTM [33] and Diffusion Policy [32] trained with 100 and 5 demonstrations. In contrast, TAPAS-GMM is only trained on five demonstrations. To experimentally disentangle policy learning and representation learning, we use groundtruth object poses as inputs for the baselines and as candidate task parameters for TAPAS-GMM. We evaluate our proposed visual features in Sec. V-C and Sec. V-D. We further compare against 3D Diffuser Actor [31], which at the time of writing, is the incumbent on RLBench. Note that it leverages a 3D scene representation, not object poses. We evaluate all policies for 200 episodes. Tab. I shows that TAPAS-GMM matches or outperforms the SOTA across all tasks while requiring significantly fewer demonstrations. Moreover, in contrast to all baselines, TAPAS-GMM scales to more complex long-horizon tasks such as SweepToDustpan and PlaceCups, without loss in performance. Moreover, unlike the Diffusion Policy and LSTM, it excels when high precision is required, e.g., in PushButton. Tab. III reports training and inference times. All GMMs are fitted on a CPU. The deep learning baselines are trained on a NVIDIA RTX A6000.
@@
Variation
Task Object Instance Clutter Objects Environment
PickAndPlace 5/5 10/10 5/5
PourDrink 17/18 10/10 5/5
Figure 7: Objects and success rates for our generalization study. Left: object instances for both tasks, Center Right: clutter objects, Right: environments.
@@
We compare the effectiveness of time-driven and state-driven GMMs in Tab. IV. To simulate disturbances during policy rollout, we reset the end-effector pose back to the starting pose, around the object contact point in the trajectory. While state-based policies are harder to learn from few demonstrations, they outperform time-based policies when confronted with disturbances due to their time-invariance. Furthermore, our action-factorized policies consistently outperform the naive implementation.
@@ V-C Real-World Policy Learning
The policy success rates in Tab. V are consistent with the results in our simulated experiments and demonstrate the core advantages of our approach: 1. strong generalization from few demonstrations, 2. handling of articulated objects, and 3. scaling to complex tasks and long horizons. Consistent with our simulated ablation experiment, TP-GMMs struggle with pose accuracy and gripper timing due to locally irrelevant task parameters and poor temporal alignment. Moreover, the deep learning baselines fail to generalize from five demonstrations and do not learn to solve any task instance. In contrast, TAPAS-GMM effortlessly scales to long-horizon tasks, without any loss in accuracy or generalization performance.
@@ V-D Visual Generalization and Skill Reuse
Prior work has extensively demonstrated the generalization capabilities of keypoints as visual features [4], which our policies inherit through the parameter generation proposed in Sec. IV-F1. We study the resulting zero-shot generalization capabilities of our policies on PickAndPlace and PourDrink by evaluating the policy on a set of unseen objects instances, clutter objects, and task environments. TAPAS-GMM exhibits strong generalization performance across the board, as shown in Fig. 7. We further provide qualitative results in the supplementary video.
Another advantage of TAPAS-GMM is that it learns multiple skill models, which can be recombined to solve novel tasks. In the supplementary video, we demonstrate how the segmented skills can be sequenced and reused in novel ways. In particular, the BakeBread task consists of taking the tray out of the oven and placing the bread on it. As an example, we invert and reorder these skills to construct a policy that takes the bread from the tray and inserts the tray into the oven.
@@ VI Conclusion
We present TAPAS-GMM, a method for learning complex manipulation tasks from only five demonstrations. By segmenting skills and automatically selecting relevant task parameters per skill, our method scales both tasks with long horizons and many parameters. Through leveraging visual semantic features, our method further generalizes across object instances, environments, and clutter. Finally, our proposed velocity factorization makes time-invariant task-parameterized policies more effective.
Limitations: TAPAS-GMM does not reason about kinematic limits or object collisions. Also, achieving decent task space coverage from only five demonstrations can be challenging with multimodal trajectory distributions. This is evident in the PlaceCups task, where workspace limits lead to vastly different trajectory shapes across task instances.